Please find I for #1. find V for #2.  find R for #3. find I, I1, I2 for #4. find R (v=12V) , I1, I2 for #5. Show work please, Thanks!4r find I I 8.52 find V 20 HE R a.. 8. find B ZA 4V 4.3.62 find I, 260 52 find 202 I, 2 lor
Answer #1

2 ร, u2 4) D H 3.60 5) R 11 w २०2 + I GV . та 6V UR 12V + R lov Reg - 3.6+ 4x6 6+4 I A = 1.5A = 3.6+ 2.4 I =15 A =62 I 12V or

Thanks for asking.